OpenStudy (anonymous):

Please help! Divide and simplify: 4x 2x ------ / ----- x^2+2x x^2+2x

OpenStudy (anonymous):

2

OpenStudy (anonymous):

"invert and multiply" \[\frac{4x}{x^2+2x}\times \frac{x^2+2x}{2x}\]

OpenStudy (anonymous):

then cancel and you will get 2